# HiCache: Training-free Acceleration of Diffusion Models via Hermite Polynomial-based Feature Caching

### v1

@@ A.4 Detailed Experimental Setup
A.4.1 Model and Task Specifications
@@
We employ the FLUX.1-dev model for high-resolution text-to-image synthesis. Images are generated at 1024×1024 resolution using 1,632 high-quality prompts sourced from the DrawBench benchmark. Quality assessment is performed using ImageReward, a robust perceptual metric for text-to-image alignment.
@@
A.4.2 Hardware and Computational Resources
@@ A.4.3 Evaluation Protocols
A.5 Super-Resolution Task: Detailed Experimental Results
@@
Method Interval
Latency
Time
Speedup
FLOPs
FLOPs
Speedup
Restoration Track
[…]
HiCache
@@
A.6 Theoretical Analysis of Gaussian Feature Dynamics and Hermite Optimality
@@
A.6.1 On the Gaussianity of Feature Derivatives (Observation 1 in the main paper)
While a complete theoretical proof of Gaussianity in neural network feature dynamics remains challenging, we provide several theoretical perspectives that support this empirical observation:
@@ A.6.1 On the Gaussianity of Feature Derivatives (Observation 1 in the main paper)
A.6.2 Hermite Polynomials as Optimal Basis (Corollary 1 in the main paper)
@@
A.6.3 Robustness Under Approximate Gaussianity
@@
A.6.4 Summary and Practical Implications
@@
1. **Theoretical Support**: Multiple theoretical frameworks (CLT, random matrix theory, diffusion approximations) suggest Gaussianity is a reasonable expectation in high-dimensional neural networks.
2. **Conditional Optimality**: Given the empirically validated Gaussian nature, the optimality of Hermite polynomials follows rigorously from the Karhunen-Loève theorem.
3. **Robustness**: Even under approximate Gaussianity, Hermite bases maintain advantages over monomials due to their oscillatory nature and better numerical conditioning.
@@
A.7 Implementation Details
This section provides the complete Python implementation of the HiCache algorithm, including all core functions for derivative approximation, feature prediction, and cache management.
1from typing import Dict, List, Union
2import torch
3import math
4
5def _hermite_polynomial(x: torch.Tensor, n: int) -> torch.Tensor:
6 """
7 Physicist’s Hermite polynomial H_n(x).
@@ A.7 Implementation Details
9 """
10 if n == 0:
11 return torch.ones_like(x)
12 elif n == 1:
13 return 2 * x
14
15 H_prev = torch.ones_like(x)
16 H_curr = 2 * x
17
18 for k in range(2, n + 1):
19 H_next = 2 * x * H_curr - 2 * (k - 1) * H_prev
20 H_prev, H_curr = H_curr, H_next
21
22 return H_curr
23
24def _collect_trajectory_feature(
25 cache_dic: Dict,
26 current: Dict,
27 feature: torch.Tensor
28):
29 """
30 Feature trajectory collector - automatically collects features during
31 cache updates. Supports simultaneous collection from multiple layers
32 and modules (collects from all required modules in a single inference
33 pass).
34
35 :param cache_dic: Cache dictionary
36 :param current: Information of the current step
37 :param feature: Feature tensor to collect
38 """
39 config = cache_dic.get(’feature_collection_config’, {})
40
41 # Support multi-layer feature collection
42 target_layers = config.get(’target_layers’, [14])
43 if isinstance(target_layers, int):
44 target_layers = [target_layers] # For backward compatibility
45
46 # Check collection conditions - collect only if in the target layer list
47 if current[’layer’] not in target_layers:
48 return
49
50 # NEW: Support simultaneous collection from multiple modules
51 target_modules = config.get(’target_modules’, [’any’])
52 if isinstance(target_modules, str):
53 target_modules = [target_modules]
54
55 # Check if the module needs to be collected
56 if ’any’ not in target_modules and current[’module’] not in target_modules:
57 return
58
59 # Support multi-stream collection
60 target_streams = config.get(’target_streams’, [’any’])
61 if isinstance(target_streams, str):
62 target_streams = [target_streams]
63
64 # Check if the stream needs to be collected
65 if ’any’ not in target_streams and current[’stream’] not in target_streams:
66 return
67
68 # NEW: Store by layer-module combination, supporting simultaneous
69 # collection from multiple modules
70 if ’trajectory_features’ not in cache_dic:
71 cache_dic[’trajectory_features’] = {}
72 cache_dic[’trajectory_metadata’] = {}
73
74 layer_key = current[’layer’]
75 module_key = current[’module’]
76
77 # Initialize layer-level storage
78 if layer_key not in cache_dic[’trajectory_features’]:
79 cache_dic[’trajectory_features’][layer_key] = {}
80 cache_dic[’trajectory_metadata’][layer_key] = {}
81
82 # Initialize module-level storage
83 if module_key not in cache_dic[’trajectory_features’][layer_key]:
84 cache_dic[’trajectory_features’][layer_key][module_key] = []
85 cache_dic[’trajectory_metadata’][layer_key][module_key] = []
86
87 # Collect features into the corresponding module storage
88 cache_dic[’trajectory_features’][layer_key][module_key].append(
89 feature.clone().detach().cpu()
90 )
91 cache_dic[’trajectory_metadata’][layer_key][module_key].append({
92 ’step’: current[’step’],
93 ’timestep’: current.get(’t’, 0),
94 ’cache_type’: current.get(’type’, ’full’),
95 ’layer’: current[’layer’],
96 ’module’: current[’module’],
97 ’stream’: current[’stream’]
98 })
99
100def derivative_approximation(
101 cache_dic: Dict,
102 current: Dict,
103 feature: torch.Tensor
104):
105 """
106 Compute derivative approximation.
107
108 :param cache_dic: Cache dictionary
109 :param current: Information of the current step
110 """
111 # NEW: Feature collection hook
112 if cache_dic.get(’enable_feature_collection’, False):
113 _collect_trajectory_feature(cache_dic, current, feature)
114 # If only for feature collection, skip the rest of the cache operations
115 if not cache_dic.get(’taylor_cache’, False):
116 return
117
118 # Safety check: ensure cache structure exists
119 try:
120 cache_path = cache_dic[’cache’][-1][current[’stream’]][current[’layer’]]
121 cache_module = cache_path[current[’module’]]
122 except KeyError:
123 return
124
125 # Safety check: ensure there are enough activated_steps to calculate the difference
126 if len(current[’activated_steps’]) < 2:
127 # If there is only one or no activated_steps, the difference cannot be calculated, so return directly
128 return
129
130 activated_steps = current[’activated_steps’]
131 difference_distance = activated_steps[-1] - activated_steps[-2]
132 #activated_times = current[’activated_times’]
133 #difference_distance = activated_times[-1] - activated_times[-2]
134
135 updated_taylor_factors = {}
136 updated_taylor_factors[0] = feature
137
138 for i in range(cache_dic[’max_order’]):
139 has_cache = cache_module.get(i, None) is not None
140 step_condition = current[’step’] > cache_dic[’first_enhance’] - 2
141 if has_cache and step_condition:
142 numerator = updated_taylor_factors[i] - cache_module[i]
143 updated_taylor_factors[i + 1] = numerator / difference_distance
144 else:
145 break
146
147 cache_path[current[’module’]] = updated_taylor_factors
148
149def get_collected_features(cache_dic: Dict) -> tuple:
150 """
151 Get the collected feature trajectories - supports multi-layer data.
152
153 :param cache_dic: Cache dictionary
154 :return: (features_dict, metadata_dict) tuple where keys are layer indices
155 """
156 features = cache_dic.get(’trajectory_features’, {})
157 metadata = cache_dic.get(’trajectory_metadata’, {})
158 return features, metadata
159
160def clear_collected_features(cache_dic: Dict):
161 """
162 Clear the collected feature trajectories.
163
164 :param cache_dic: Cache dictionary
165 """
166 if ’trajectory_features’ in cache_dic:
167 del cache_dic[’trajectory_features’]
168 if ’trajectory_metadata’ in cache_dic:
169 del cache_dic[’trajectory_metadata’]
170
171
172def taylor_formula(cache_dic: Dict, current: Dict) -> torch.Tensor:
173 """
174 Feature prediction dispatcher: chooses between Taylor or Hermite prediction.
175
176 :param cache_dic: Cache dictionary
177 - ’prediction_mode’: ’taylor’ or ’hermite’. Defaults to ’taylor’ if not specified.
178 - ’use_hermite’: (Legacy) If True and ’prediction_mode’ is not set, mode becomes ’hermite’.
179 :param current: Information of the current step
180 """
181 # Determine prediction mode, with backward compatibility for ’use_hermite’
182 if ’prediction_mode’ in cache_dic:
183 mode = cache_dic[’prediction_mode’]
184 elif cache_dic.get(’use_hermite’, False):
185 mode = ’hermite’
186 else:
187 mode = ’taylor’
188
189 # Dispatch based on mode
190 if mode == ’taylor’:
191 return _taylor_expansion_formula(cache_dic, current)
192
193 elif mode == ’hermite’:
194 return _hermite_prediction_formula(cache_dic, current)
195
196 else:
197 error_msg = (f"Unknown prediction_mode: ’{mode}’. "
198 f"Must be ’taylor’ or ’hermite’.")
199 raise ValueError(error_msg)
200
201
202def _taylor_expansion_formula(cache_dic: Dict, current: Dict) -> torch.Tensor:
203 """
204 Standard Taylor expansion prediction.
205 Uses a power function basis: F_pred = F_0 + Sum (1/k!) * x^k * Delta^kF
206 """
207 activated_steps = current[’activated_steps’]
208 x = current[’step’] - activated_steps[-1]
209 #x = current[’t’] - current[’activated_times’][-1]
210 output = 0
211
212 # FIX: Safety check, ensure cache structure exists
213 try:
214 cache_root = cache_dic[’cache’][-1]
215 stream_cache = cache_root[current[’stream’]]
216 layer_cache = stream_cache[current[’layer’]]
217 feats_d = layer_cache[current[’module’]]
218 except KeyError:
219 # FIX: If the cache does not exist, it means this is the first access to this module.
220 # In this case, we should initialize the cache structure.
221 cache_root = cache_dic[’cache’][-1]
222 if current[’stream’] not in cache_root:
223 cache_root[current[’stream’]] = {}
224 stream_cache = cache_root[current[’stream’]]
225 if current[’layer’] not in stream_cache:
226 stream_cache[current[’layer’]] = {}
227 layer_cache = stream_cache[current[’layer’]]
228 if current[’module’] not in layer_cache:
229 layer_cache[current[’module’]] = {}
230
231 # FIX: If the cache is empty, it means this is the first call to this module.
232 # In this case, we should throw a more user-friendly error message,
233 # guiding the user to check the configuration.
234 error_msg = (f"Cache not found for stream=’{current[’stream’]}’, "
235 f"layer={current[’layer’]}, module=’{current[’module’]}’. "
236 f"This usually means the first step was not run in ’full’ "
237 f"mode to initialize the cache. Please check your cache "
238 f"configuration and ensure first_enhance >= 1.")
239 raise ValueError(error_msg)
240
241 # FIX: Use the max_order parameter to limit the number of terms used
242 max_order = cache_dic.get(’max_order’, 3)
243 # +1 because we include 0th order
244 effective_order = min(max_order + 1, len(feats_d))
245
246 for i in range(effective_order):
247 output += (1 / math.factorial(i)) * feats_d[i] * (x ** i)
248
249 return output
250
251
252def _hermite_prediction_formula(
253 cache_dic: Dict,
254 current: Dict
255) -> torch.Tensor:
256 """
257 Feature prediction based on Hermite polynomials.
258
259 Uses Hermite polynomials as basis functions instead of standard power
@@ 30 Feature trajectory collector - automatically collects features during
261
262 Where:
263 - H_k(x) is the k-th order Hermite polynomial
264 - Delta^kF is the k-th order difference feature
265 - x is the time step difference
266
267 Advantages of this method:
268 1. Hermite polynomials are orthogonal, providing better numerical
269 stability.
270 2. Higher approximation accuracy for certain types of functions.
271 3. The growth of the polynomial can be controlled by a scaling factor.
272 """
273 # Get the actual time step difference, consistent with the original
274 # Taylor expansion
275 activated_steps = current[’activated_steps’]
276 x = current[’step’] - activated_steps[-1]
277 # If time difference is needed, uncomment the line below
278 # x = current[’t’] - current[’activated_times’][-1]
279
280 # Get feature cache
281 try:
282 cache_root = cache_dic[’cache’][-1]
283 stream_cache = cache_root[current[’stream’]]
284 layer_cache = stream_cache[current[’layer’]]
285 feats_d = layer_cache[current[’module’]]
286 except KeyError:
287 error_msg = (f"Cache not found for stream=’{current[’stream’]}’, "
288 f"layer={current[’layer’]}, module=’{current[’module’]}’")
289 raise ValueError(error_msg)
290
291
292 # FIX: Use the max_order parameter to limit the order
293 max_order = cache_dic.get(’max_order’, 3)
294 # available order = number of historical terms - 1
295 available_order = len(feats_d) - 1
296 # use the smaller value
297 order = min(max_order, available_order)
298
299 if order < 1:
300 # Not enough history, return the latest feature
301 return feats_d.get(0)
302
303 # F_0
304 F_latest = feats_d[0].clone()
305
306 # Convert the time step difference to a tensor, keeping the same
307 # dtype and device as the feature
308 x_tensor = torch.tensor(
309 float(x),
310 dtype=F_latest.dtype,
311 device=F_latest.device
312 )
313
314 # Get the scaling factor to control the growth of the Hermite polynomial
315 scale_factor = cache_dic.get(’hermite_scale_factor’, 0.5)
316 x_scaled = x_tensor * scale_factor
317
318 # Construct Hermite prediction
319 pred = F_latest.clone()
320
321 for k in range(1, order + 1):
322 diff_k = feats_d[k]
323 Hk = _hermite_polynomial(x_scaled, k)
324 # Consider the effect of the scaling factor
325 alpha = float(Hk / math.factorial(k)) * (scale_factor ** k)
326 pred.add_(diff_k, alpha=alpha)
327
328 return pred
329
330
331
332def taylor_cache_init(cache_dic: Dict, current: Dict):
333 """
334 Initialize Taylor cache and allocate storage for different-order
335 derivatives in the Taylor cache.
336
337 :param cache_dic: Cache dictionary
338 :param current: Information of the current step
339 """
340 if (current[’step’] == 0) and (cache_dic[’taylor_cache’]):
341 cache_root = cache_dic[’cache’][-1]
342 stream_cache = cache_root[current[’stream’]]
343 layer_cache = stream_cache[current[’layer’]]
344 layer_cache[current[’module’]] = {}
Listing 1: Core Implementation of HiCache Algorithm
The implementation includes several key components:
  • Hermite Polynomial Computation (_hermite_polynomial): Recursive computation using the recurrence relation in Eq. (14)
  • Feature Collection (_collect_trajectory_feature): Multi-layer, multi-module feature trajectory collection
  • Derivative Approximation (derivative_approximation): Finite difference approximations as described in Algorithm 1
  • Prediction Dispatcher (taylor_formula): Interface supporting both Taylor and Hermite prediction modes
  • Hermite Prediction (_hermite_prediction_formula): Hermite polynomial-based feature prediction as described in Algorithm 2
  • Taylor Prediction (_taylor_expansion_formula): Taylor expansion implementation

### v2

@@ Appendix E Detailed Experimental Setup
Model and Task Specifications
@@
We employ the FLUX.1-dev model for high-resolution text-to-image synthesis. Images are generated at 1024x1024 resolution using 200 high-quality prompts sourced from the DrawBench benchmark. Quality assessment is performed using ImageReward, a robust perceptual metric for text-to-image alignment.
@@
Hardware and Computational Resources
@@
Appendix F Super-Resolution Task: Detailed Experimental Results
@@
Method Interval Latency (s) Time Speedup FLOPs (G) FLOPs Speedup Restoration Track
@@
Appendix G Theoretical Analysis of Gaussian Feature Dynamics and Hermite Optimality
@@
On the Gaussianity of Feature Derivatives
We present two complementary, verifiable mechanisms that yield quantitative Gaussian approximations for finite differences $\Delta\mathbf{F}_{t}$.
Proposition 4 (Local linearization $\Rightarrow$ approximate Gaussian).
Let $x_{t}=\alpha_{t}x_{0}+\sigma_{t}\,\varepsilon_{t}$ with $\varepsilon_{t}\!\sim\!\mathcal{N}(0,I)$ independent of $x_{0}$. Let $F:\mathbb{R}^{d}\!\times\!\mathbb{R}\to\mathbb{R}^{p}$ be twice differentiable with bounded second derivatives $\|\nabla_{x}^{2}F\|_{\text{op}}\leq L_{x}$, $|\partial_{tt}F|\leq L_{t}$, $\|\partial_{t}\nabla_{x}F\|_{\text{op}}\leq L_{xt}$ on a neighborhood. For a small step $\delta>0$, define $\Delta F_{t}:=F(x_{t},t)-F(x_{t-\delta},t-\delta)$. Then there exist $A_{t},b_{t}$ and residual $R_{t}$ such that
$\Delta F_{t}=A_{t}\,(\varepsilon_{t}-\tilde{\varepsilon}_{t-\delta})+b_{t}+R_{t},\qquad\tilde{\varepsilon}_{t-\delta}\!\sim\!\mathcal{N}(0,I),$ (27)
with $\mathbb{E}\,\|R_{t}\|_{2}\leq C\,(L_{x}\,\mathbb{E}\|x_{t}-x_{t-\delta}\|_{2}^{2}+L_{t}\,\delta^{2}+L_{xt}\,\delta\,\mathbb{E}\|x_{t}-x_{t-\delta}\|_{2})=O(\delta)$.
Proof.
By the two-variable Taylor formula around $(x_{t},t)$,
$F(x_{t},t)-F(x_{t-\delta},t-\delta)=\nabla_{x}F(x_{t},t)(x_{t}-x_{t-\delta})-\partial_{t}F(x_{t},t)\,\delta+R_{t},$ (28)
with $\mathbb{E}\,\|R_{t}\|_{2}$ controlled by the stated bounds. Using $x_{t}-x_{t-\delta}=(\alpha_{t}-\alpha_{t-\delta})x_{0}+\sigma_{t}\varepsilon_{t}-\sigma_{t-\delta}\varepsilon_{t-\delta}$ and grouping constants into $b_{t}$ yields the claimed linear-Gaussian leading term. ∎
Proposition 5 (Aggregation $\Rightarrow$ CLT with Berry–Esseen bound).
Suppose $\Delta F_{t}=\sum_{i=1}^{M}Y_{i,t}$ where $Y_{i,t}\in\mathbb{R}^{p}$ are zero-mean contributions (e.g., heads/tokens/neurons) with uniformly bounded third moments and a bounded-degree dependency graph. Let $S_{t}=\sum_{i}\mathrm{Cov}(Y_{i,t})$ and define $\eta_{t}:=\max_{\|u\|=1}\frac{\sum_{i}\mathbb{E}|u^{\top}Y_{i,t}|^{3}}{(u^{\top}S_{t}u)^{3/2}}$. Then for any unit $u$,
$\sup_{x\in\mathbb{R}}\Big|\mathbb{P}\big(u^{\top}\!\Delta F_{t}\leq x\big)-\Phi\!\Big(\frac{x}{\sqrt{u^{\top}S_{t}u}}\Big)\Big|\;\leq\;C\,\eta_{t},$ (29)
and $\Delta F_{t}\Rightarrow\mathcal{N}(0,S_{t})$ as $\eta_{t}\to 0$.
Proof.
By the Cramér–Wold device, analyze $u^{\top}\Delta F_{t}=\sum_{i}X_{i}$ with $X_{i}:=u^{\top}Y_{i,t}$. A Berry–Esseen bound for bounded-degree dependency graphs controls the Kolmogorov distance by $C\,\eta_{t}$. Uniformity over $u$ yields the multivariate statement. ∎
Diagnostics.
We monitor Gaussianity via energy tests, dispersion proxies for $\eta_{t}$ (e.g., effective number of contributors), and local linearization error versus $\delta$.
@@
Hermite Polynomials as Optimal Basis
@@
Proposition 7 (Projection optimality in $L^{2}(\gamma)$).
Let $\gamma$ be a Gaussian measure on the horizon variable. For any target offset function $y$ and any degree-$N$ polynomial $p$, the orthogonal projection $P_{N}y$ of $y$ onto $\mathrm{span}\{H_{0},\ldots,H_{N}\}$ satisfies
$\|y-P_{N}y\|_{2,\gamma}\leq\|y-p\|_{2,\gamma}.$ (35)
Thus truncated Hermite expansion minimizes the Gaussian-weighted L2 error within degree-$N$ polynomials.
Proof.
Hermite polynomials form an orthogonal basis of $L^{2}(\gamma)$. For any closed subspace $\mathcal{V}_{N}:=\mathrm{span}\{H_{0},\ldots,H_{N}\}$, the orthogonal projection $P_{N}$ minimizes distance to $\mathcal{V}_{N}$ by the Pythagorean theorem: for any $v\in\mathcal{V}_{N}$, $\|y-P_{N}y\|_{2,\gamma}\leq\|y-v\|_{2,\gamma}$. Choosing $v=p$ gives the claim. ∎
@@
Robustness Under Approximate Gaussianity
@@
Summary and Practical Implications
@@
  1. 0.1
    Theoretical Support: Multiple theoretical frameworks (CLT, random matrix theory, diffusion approximations) suggest Gaussianity is a reasonable expectation in high-dimensional neural networks.
  2. 0.2
    Conditional Optimality: Given the empirically validated Gaussian nature, the optimality of Hermite polynomials follows rigorously from the Karhunen-Loève theorem.
  3. 0.3
    Robustness: Even under approximate Gaussianity, Hermite bases maintain advantages over monomials due to their oscillatory nature and better numerical conditioning.
@@
Appendix H Additional Qualitative Results
Empirical Verification of Oscillatory Advantages
Figure 9 provides empirical evidence for the theoretical advantages of Hermite polynomials’ oscillatory properties discussed in the main paper. Using the non-cumulative error framework on representative FLUX features, we observe that while both methods perform similarly at order 1, Taylor expansion increasingly over-extrapolates at higher orders, especially at trajectory turning points—precisely confirming our theoretical analysis about monomial limitations. HiCache’s oscillatory Hermite basis, stabilized by the contraction factor $\sigma$, maintains accurate fitting across all orders.
Appendix I Additional Experimental Results
This section reports additional experiments that complement the main text: (i) empirical Gaussianity tests for finite differences on FLUX.1-dev, (ii) an adaptive per-layer $\sigma$ scheme motivated by our truncation-error analysis, and (iii) extra backbones and accelerators (Qwen-Image and Chipmunk-Flux).
Empirical Gaussianity via Energy Distance Tests on FLUX.1-dev
To support the Gaussian finite-difference assumption underlying our Hermite basis, we conduct a comprehensive energy-distance test on FLUX.1-dev features. We collect internal trajectories at six representative layers (4/10/14/20/28/42), covering both early dual-stream and late single-stream blocks, and separately track image/text attention and MLP streams when applicable. For each (layer, module) configuration we compute 1st–5th order finite differences over timesteps and apply a high-dimensional energy distance test (Székely & Rizzo) against a multivariate Gaussian fitted by the empirical mean and covariance, using 500 parametric bootstrap simulations and 400 randomly sampled spatial patches.
Using all timesteps, we obtain $15\times 5=75$ tests in total and observe that all p-values are above $0.05$. We further split the trajectory into three regimes on the finite-difference time axis: early high-noise ([0,15)), mid transition ([15,35)), and late low-noise ([35,50)), and rerun the tests in each regime (again 75 tests per regime). The aggregated pass rates are summarized in Table 6.
Table 6: Energy-distance Gaussianity tests on FLUX.1-dev finite differences (15 layer–module configurations $\times$ 5 orders). “Pass” denotes p-value >0.05>0.05.
@@ Appendix I Additional Experimental Results
All timesteps (all layers/modules) $15\times 5$ 75 75 0 100%
Early high-noise phase $15\times 5$ 75 75 0 100%
Mid transition phase $15\times 5$ 75 73 2 97.3%
Late low-noise phase $15\times 5$ 75 75 0 100%
Overall, finite-difference features on FLUX.1-dev are empirically very close to Gaussian across depth, modality, and time: all 75 all-timestep tests and 150/150 tests in the early/late phases pass at the 0.05 level, and only 2/75 high-order shallow-text configurations in the mid-transition phase fall slightly below 0.05. This supports the modeling assumption that motivates our Hermite basis.
Adaptive Per-Layer $\sigma$ on FLUX.1-dev
Section 3 analyzes Hermite truncation errors and shows that stable extrapolation requires a condition of the form $\sigma\sqrt{2}\lvert\Delta s\rvert\lesssim 1$. Rather than manually tuning a single global $\sigma$ (e.g., $\sigma=0.5$), we design a training-free adaptive scheme that chooses a per-layer $\sigma_{\ell}$ based on the observed scale of feature differences, while respecting this stability constraint.
For each layer $\ell$, we track first-order finite differences $\Delta F_{\ell,t}$ and compress them into a scalar magnitude $D_{\ell,t}$ via an $\ell_{2}$ norm over channels followed by averaging over space and batch. We then maintain an EMA-style RMS estimate
$q_{\ell}^{(t)}=\sqrt{(1-\beta)\,\bigl(q_{\ell}^{(t-1)}\bigr)^{2}+\beta\,D_{\ell,t}^{2}},$ (41)
initialized at $q_{\ell}^{(0)}\approx 1$. The per-layer contraction factor is set to
$\sigma_{\ell}=\min\!\left(\sigma_{\max},\frac{\alpha}{\sqrt{2}\,q_{\ell}+\varepsilon}\right),$ (42)
where $\alpha>0$ determines the desired initial scale (e.g., $\sigma_{\ell}\approx 0.5$ or $\approx 1.0$ when $q_{\ell}\approx 1$), $\sigma_{\max}$ caps large values for stability, and $\varepsilon>0$ avoids division by zero. In practice this adaptive rule simply replaces the fixed scale factor in HiCache with $\sigma_{\ell}$, without adding any learned parameters.
We evaluate this adaptive $\sigma$ on FLUX.1-dev ($\times1024$, 50 steps, interval =7=7, order =2=2, 200 prompts) and compare against TaylorSeer and fixed-$\sigma$ HiCache. All configurations share the same acceleration setting (interval =7=7, max_order =2=2). Table 7 reports CLIP, PSNR, SSIM, LPIPS, and ImageReward.
Table 7: Adaptive per-layer $\sigma$ (HiCache-Adaptive) on FLUX.1-dev under interval =7=7, order =2=2 (200 prompts, $\times1024$). Here $\sigma_{\text{fixed}/\max}$ denotes the global $\sigma$ for fixed HiCache and $\sigma_{\max}$ for HiCache-Adaptive.
@@
Baseline
TaylorSeer 27.41 28.63 0.621 0.456 0.951
Comparison at $\sigma\approx 1.0$
HiCache-Fixed 1.0 27.19 28.10 0.362 0.725 0.736
HiCache-Adaptive 1.40 1.0 $\approx 1.0$ 27.80 29.14 0.648 0.415 0.957
Comparison at $\sigma\approx 0.5$
HiCache-Fixed 0.5 27.62 28.93 0.655 0.404 0.974
HiCache-Adaptive 0.70 0.7 $\approx 0.5$ 27.87 29.11 0.646 0.415 0.971
In the moderate regime where fixed $\sigma=0.5$ already provides strong performance, adaptive $\sigma$ closely matches or slightly improves this baseline (differences in the reported metrics are within $0.02$). When $\sigma$ is increased to $1.0$, the fixed-$\sigma$ variant becomes unstable (SSIM decreases from $\approx 0.655$ to $\approx 0.362$ and LPIPS increases from $\approx 0.404$ to $\approx 0.725$), whereas adaptive $\sigma$ with initial scale $\approx 1.0$ remains stable and yields quality close to the $\sigma=0.5$ regime. This supports the theoretical motivation that the adaptive update can automatically prevent overly aggressive $\sigma$ choices while preserving performance when the baseline configuration is already well behaved.
Qwen-Image Text-to-Image Experiments
We next evaluate HiCache on Qwen-Image [Wu et al., 2025] to assess whether the conclusions from FLUX transfer to a different text-to-image backbone. All runs use the same diffusion sampler and hyperparameters; the only differences are whether TaylorSeer or HiCache is enabled and the acceleration interval. We report CLIP, ImageReward, PSNR, SSIM, and LPIPS over 200 prompts at $\times1328$ resolution with 50 sampling steps.
Table 8 summarizes the results for intervals $3/6/7/8$. At interval 3 (mild acceleration) TaylorSeer and HiCache obtain very similar scores. For larger intervals (6–8), HiCache consistently achieves higher CLIP, ImageReward, and lower LPIPS than TaylorSeer, while keeping PSNR and SSIM at a comparable level, indicating that the Hermite-based forecast is more robust under stronger acceleration.
Table 8: Qwen-Image results ($\times1328$, 50 steps, 200 prompts) comparing TaylorSeer and HiCache at different intervals, with approximate FLOPs speedup.
[…]
3 Taylor 2.78 28.96 1.214 30.44 0.800 0.208
HiCache 2.78 28.99 1.216 30.92 0.795 0.198
6 Taylor 5.00 28.09 1.012 28.52 0.601 0.481
HiCache 5.00 28.62 1.070 28.72 0.613 0.422
7 Taylor 5.56 27.73 0.895 28.34 0.564 0.538
HiCache 5.56 27.87 0.944 28.50 0.579 0.478
8 Taylor 6.25 27.01 0.750 28.23 0.519 0.591
HiCache 6.25 27.72 0.844 28.37 0.539 0.531
Chipmunk-Flux: Compatibility with Column-Sparse Accelerators
Finally, we examine HiCache on top of Chipmunk-Flux, a column-sparse accelerator for FLUX, to verify that HiCache is complementary to sparse/efficient attention rather than a competing alternative. We compare three configurations on flux-dev ($\times1024$, 50 steps, 200 prompts): pure Chipmunk-Flux, Chipmunk-Flux+Taylor, and Chipmunk-Flux+HiCache with a single $\sigma$ setting. All accelerated variants share the same interval =5=5 and order =2=2.
Table 9 reports CLIP, ImageReward, PSNR, SSIM, LPIPS, and approximate per-image latency. Moving from pure Chipmunk-Flux to Chipmunk+Taylor significantly reduces latency but degrades quality, especially in ImageReward and LPIPS. Chipmunk+HiCache with $\sigma=0.5$ largely recovers the quality of the pure Chipmunk baseline—often slightly exceeding it in ImageReward—while retaining latency close to Chipmunk+Taylor. This suggests that HiCache can serve as a plug-in quality-restoring module on top of column-sparse accelerators.
Table 9: Chipmunk-Flux experiments (flux-dev, $\times1024$, 50 steps, 200 prompts). All accelerated configurations use interval =5=5, order =2=2.
[…]
Pure (Chipmunk-Flux) 27.73 0.931 28.22 0.434 0.677 5.8
+ Taylor 27.34 0.845 28.12 0.380 0.746 3.3
+ HiCache ($\sigma{=}0.5$) 27.88 0.938 28.12 0.415 0.704 3.5